# LocoMan: Advancing Versatile Quadrupedal Dexterity with Lightweight Loco-Manipulators

### v1

@@ V-E1 Manipulation with 6DoF End Effector
To validate that LocoMan is capable of performing manipulations in the 6D operational space, we evaluate LocoMan on two challenging tasks: opening a sliding drawer and the swing door of a cabinet. These two tasks demand precise control over the robot end effector pose throughout the interaction. As illustrated in Fig. 7 (a), when opening the sliding drawer, LocoMan successfully controls the end effector to move linearly along the sliding direction of the drawer while maintaining a stable orientation axis parallel to the movement path. Fig. 7(b) demonstrates LocoMan’s capability in opening a cabinet swing door, where it maneuvers the end effector in an arc trajectory and adjusts the gripper orientation to stay perpendicular to the door plane to avoid collision with the door handle.
Furthermore, we teleoperate LocoMan to perform two additional tasks that require precise motion control: inserting a charger into a power socket and pouring liquid from one cup to another. As illustrated in Fig. 8 (a), LocoMan can accurately align the charger with the socket, demonstrating its ability to precisely adjust its end effector position. In Fig. 8 (b), LocoMan pours liquid from one cup to another, which highlights the robot’s ability to stably modulate the end effector orientation.
@@ V-E2 Manipulation in Narrow Space
Fetching objects from extremely narrow space is a challenge for quadrupedal robots due to the limited space for adjusting both the leg and manipulator joint angles. As illustrated in Fig. 9 (a), the task requires LocoMan to walk below a bed to grasp an object located under a cabinet and then carry it out. This confined environment only offers a $0.25$ $m$ clearance below the bed and an even more constricted $0.09$ $m$ height under the cabinet. As illustrated in Fig. 9 (b), LocoMan can walk stably below the vertically restricted space, avoiding any contact with the bed. This is only possible without additional robot arm installed on the top such as the ones in [18, 21]. Thanks to the compact design of loco-manipulator, LocoMan can deftly maneuver its end effector to extend under the cabinet and adjust its pose to successfully grasp the object. LocoMan then retrieves its loco-manipulator while holding the object in the gripper and exits the narrow space.
The successful completion of this challenging task demonstrates LocoMan’s competency in loco-manipulation within narrow and complex environments. This showcases its potential for deployment in scenarios such as search and rescue missions, which often necessitate both agile locomotion and manipulation in cramped spaces.
Figure 9: LocoMan demonstrates its locomotion and manipulation capabilities in constrained environments under joystick teleoperation. (a) The task requires the robot to walk beneath a crib with a clearance of 0.25 m and grasp the object under the cabinet with a clearance of 0.09 m. (b) LocoMan crawls forward to approach the targeted object. (b) LocoMan extends its loco-manipulator to reach and grasp the object. (d) LocoMan retrieves the object. (e) LocoMan crawls backward and exits the crib.
@@ V-E3 Bimanual Manipulation
Hoisting objects with a rope is a task that demands precise bimanual coordination to ensure the rope does not slip back. As illustrated in Fig. 10 (a), while maintaining an upright stance, LocoMan adeptly adjusts the 6D pose of an end effector to securely grasp the rope and applies a downward pull. Upon reaching the limits of its workspace, the pulling end effector maintains its grip on the rope, steadfastly holding it in place until the alternate end effector takes over to continue pulling the rope. The seamless transition and successful execution of the hoisting operation vividly illustrate LocoMan’s advanced bimanual coordination capabilities, which showcase the practical versatility and enhanced operational scope of quadrupedal robots equipped with our proposed loco-manipulators.
@@ VI Conclusion
In this paper, we present LocoMan, a novel approach that enhances the manipulation dexterity of quadrupedal robots through the integration of designed lightweight loco-manipulators, expanding their operational workspace and enabling precise control over complex 6D manipulation tasks. Our design effectively combine the mobility of the quadrupedal robots with the functionality of manipulators, without compromising agility or requiring extensive payload capacity. The developed unified control framework assures accurate and stable movement across a spectrum of tasks, illustrating LocoMan’s versatility in environments ranging from confined spaces to tasks requiring intricate dual-arm coordination. In the future, we would utilize this platform to train a wider range of quadrupedal loco-manipulation tasks through imitation learning, and achieve smooth and efficient mode switch by reinforcement learning.
@@
VI-A Joint Limits of Loco-Manipulator
@@
VI-B Whole-Body Impulse Control
@@
VI-B1 Hierarchical Objective Tracking
@@ VI-B2 Quadratic Programming
where the optimization objective is to minimize the reaction force $\mathbf{f}_{r}$ and the relaxation variable $\bm{\delta}_{t}$ for the floating torso acceleration. $\bm{S}_{f}$, $\bm{A}$, $\mathbf{b}$, $\mathbf{g}$, and $\bm{W}$ are the floating torso selection matrix, the mass matrix, the Coriolis force, the gravitation force, and the force constraint matrix for foot contact, respectively.
@@
VI-B3 Joint-Level PD Control
@@
VI-B4 PD Gains for Real Robot

### v2

@@ V-E1 Manipulation with 6DoF End Effector
To validate that LocoMan is capable of performing manipulations in the 6D operational space, we evaluate LocoMan on two challenging tasks: opening a sliding drawer and the swing door of a cabinet. These two tasks demand precise control over the robot end effector pose throughout the interaction.
As illustrated in Fig. 7 (a), when opening the sliding drawer, LocoMan successfully controls the end effector to move linearly along the sliding direction of the drawer while maintaining a stable orientation axis parallel to the movement path. Fig. 7(b) demonstrates LocoMan’s capability in opening a cabinet swing door, where it maneuvers the end effector in an arc trajectory and adjusts the gripper orientation to stay perpendicular to the door plane to avoid collision with the door handle.
Furthermore, we teleoperate LocoMan to perform two additional tasks that require precise motion control: inserting a charger into a power socket shown in Fig. 8 (a), LocoMan can accurately align the charger with the socket, demonstrating its ability to precisely adjust its end effector position. In Fig. 8 (b), LocoMan pours liquid from one cup to another, which highlights its ability to stably modulate the end effector orientation.
@@ V-E2 Manipulation in Narrow Space
Fetching objects from an extremely narrow space is a challenge for quadrupedal robots due to the limited space for adjusting both the leg and manipulator joint angles.
As illustrated in Fig. 9 (a), the task requires LocoMan to walk below a bed to grasp an object located under a cabinet and then carry it out. This confined environment only offers a $0.25$ $m$ clearance below the bed and an even more constricted $0.09$ $m$ height under the cabinet. As illustrated in Fig. 9 (b), LocoMan can walk stably below the vertically restricted space, avoiding any contact with the bed. This is only possible without an additional robot arm installed on the top, such as those in [18, 21]. Thanks to the compact design of loco-manipulator, LocoMan can deftly maneuver its end effector to extend under the cabinet and adjust its pose to successfully grasp the object. LocoMan then retrieves its loco-manipulator while holding the object in the gripper and exits the narrow space. The successful completion of this challenging task demonstrates LocoMan’s competency in loco-manipulation within narrow and complex environments. This showcases its potential for deployment in scenarios such as search and rescue missions, which often necessitate both agile locomotion and manipulation in cramped spaces.
Figure 9: LocoMan demonstrates its locomotion and manipulation capabilities in constrained environments under joystick teleoperation. (a) The task requires the robot to walk beneath a crib with a clearance of 0.25 m and grasp the object under the cabinet with a clearance of 0.09 m. (b) LocoMan crawls forward to approach the targeted object. (c) LocoMan extends its loco-manipulator to reach and grasp the object. (d) LocoMan retrieves the object. (e) LocoMan crawls backward and exits the crib.
@@ V-E3 Bimanual Manipulation
Hoisting objects with a rope is a task that requires precise bimanual coordination to ensure the rope does not slip back. As illustrated in Fig. 10 (a), while maintaining an upright stance, LocoMan adeptly adjusts the 6D pose of an end effector to securely grasp the rope and applies a downward pull. Upon reaching the limits of its workspace, the pulling end effector maintains its grip on the rope, steadfastly holding it in place until the alternate end effector takes over to continue pulling the rope. The seamless transition and successful execution of the hoisting operation vividly illustrate LocoMan’s advanced bimanual coordination capabilities, which showcase the practical versatility and enhanced operational scope of quadrupedal robots equipped with our proposed loco-manipulators.
@@ VI Conclusion
In this paper, we present LocoMan, a novel approach that enhances the manipulation dexterity of quadrupedal robots through the integration of two lightweight loco-manipulators, expanding their operational workspace and enabling precise control over complex 6D manipulation tasks. Our design effectively combines the mobility of quadrupedal robots with the functionality of manipulators, without compromising agility or requiring extensive payload capacity. The developed unified control framework ensures accurate and stable movement across a spectrum of tasks, illustrating LocoMan’s versatility in environments ranging from confined spaces to tasks requiring intricate dual-arm coordination.
VII Acknowledgements
The work is partially supported by Google Deepmind with an unrestricted grant. The authors also acknowledge support from the National Science Foundation under grants CNS-2047454.
@@ VII Acknowledgements
VII-A Joint Limits of Loco-Manipulator
@@
VII-B Whole-Body Impulse Control
@@
VII-B1 Hierarchical Objective Tracking
@@ VII-B2 Quadratic Programming
where the optimization objective is to minimize the reaction force $\mathbf{f}_{r}$ and the relaxation variable $\bm{\delta}_{t}$ for the floating torso acceleration with the weights $\bm{Q}_{1}$ and $\bm{Q}_{2}$. $\bm{S}_{f}$, $\bm{A}$, $\mathbf{b}$, $\mathbf{g}$, and $\bm{W}$ are the floating torso selection matrix, the mass matrix, the Coriolis force, the gravitation force, and the force constraint matrix for foot contact, respectively.
@@
VII-B3 Joint-Level PD Control
@@
VII-B4 PD Gains for Real Robot